OpenStudy (anonymous):

can anybody help me with the connexus algebra 2 lesson 11 unit 4?

OpenStudy (amistre64):

if you post a question, it would be easier to help ....

OpenStudy (anonymous):

a football is kicked into the air from an initial height of 3 feet. the height, in feet, of the football above the ground is given by s(t)=-16t^2+35t+3... where t is time in seconds and t>=0. which is closest to the time when the football will be 20ft above the ground??

OpenStudy (anonymous):

2.72 seconds 0.73 seconds or 1.46 seconds 0.53 seconds 0.53 seconds or 2.72 seconds

OpenStudy (amistre64):

do you have any ideas on how to approach the solution?

OpenStudy (anonymous):

no i do not, im struggling BIG TIME with algebra

OpenStudy (amistre64):

one way to approach the solution will depend on knowing the quadratic formula

OpenStudy (anonymous):

i dont understand how the question is, and how to use the quadratic formula with it.

OpenStudy (amistre64):

using the quadratic formula is a simple plug and compute operation .... the question gives us a quadratic equation the defines the height at any moment in time (t). to determine when that equation is equal to 20, set it equal to 20... s(t)=-16t^2+35t+3 ; when does s(t)= 20? 20 = -16t^2+35t+3 subtract 20 from both sides: 0 = -16t^2+35t -17 now we have a format that we can use the quadratic equation with

OpenStudy (amistre64):

given a setup: \( ax^2+bx+c=0\) the quadratic formula defines the variable x as:\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):



OpenStudy (amistre64):

thats pretty good, missed a b^2 is all i see

OpenStudy (anonymous):

yea i noticed that lol

OpenStudy (anonymous):

but that confuses me still . i understand how to do it when my teacher goes over it, but then when i do it myself i dont get it

OpenStudy (amistre64):

the proof behind the quadratic formula is developed from a method called; completing the square

OpenStudy (amistre64):

\[0=ax^2+bx+c\] \[0=a(x^2+\frac bax+\frac ca)\] \[0=x^2+\frac bax+\frac ca\] \[\frac{b^2}{4a^2}-\frac ca=x^2+\frac bax+\frac{b^2}{4a^2}\] \[\frac{b^2}{4a^2}-\frac ca=(x+\frac ba)^2\] \[\pm\sqrt{\frac{b^2}{4a^2}-\frac ca}=x+\frac ba\] \[-\frac{b}{a}\pm\sqrt{\frac{b^2}{4a^2}-\frac ca}=x\] to put it into its "usual" form you can then paly with the adding the fractions and simplifying

OpenStudy (amistre64):

i dropped a 2 along the way trying to code it nice :/

OpenStudy (amistre64):

\[\frac{b^2}{4a^2}-\frac ca=(x+\frac ba)^2\] ^^ 2a \[\frac{b^2}{4a^2}-\frac ca=(x+\frac b{2a})^2\] and then it works out as expected :)

OpenStudy (anonymous):

ok. ill try it. ty for ur help .

OpenStudy (amistre64):

good luck